mix 6.80 mL of CHCl3 (d=1.492 g/mL) and 6.90 mL of CHBr3 (d= 2.890 g/mL) giving 13.7 mL



1
Expert's answer
2021-08-30T02:25:53-0400

Question:

Consider the following: you mix 6.80mL of CHCl3 (d = 1.492 g/mL) and 6.90mL of CHBr3 (d = 2.890 g/mL), giving 13.7mL of solution. What is the density of this mixture?


Explanation:

The density of mixture will be calculated by taking ratio of total mass of the mixture obtained by mixing the two liquids and the total volume of the mixture


The total volume = volume of "CHCl_3" + volume of "CHBr_3" = "6.80+6.90 =13.70mL"


The mass of CHCl3-l; = density of CHCI3 × volume of ICI3

= "1.492 × 6.80= 10.15
\ g"


the mass of CHBr3-I; = density of CHBr3 × volume of CHBrl3 "= 2.890 × 6.90= 19.94\ g"


Thus total mass "= 10.15\ g + 19.94\ g = 30.09\ g"


Density of mixture = total mass / total volume "=30.09/13.70 = 2.20\ g/mL"